Suppose the following table was generated from the sample data of 2020 employees relating annual salary to years of education and gender. According to the results, is there a salary difference between men and women at the 0.050.05 level of significance? If yes, write the difference in salary in the space provided, rounded to two decimal places. Else, select "There is not enough evidence."

Coefficients Standard Error t Stat P-Value
Intercept −8619.401339−8619.401339 3672.4312293672.431229 −2.347056−2.347056 0.0312940.031294
Education 3622.7871453622.787145 231.963588231.963588 15.61791315.617913 0.0000000.000000
Female (1 if female, 0 if male) 4905.2082944905.208294 1461.7387531461.738753 3.3557353.355735 0.0037500.003750

Selecting a radio button will replace the entered answer value(s) with the radio button value. If the radio button is not selected, the entered answer is used.
Answer #1

Regression equation:

Salary = −8619.401339 + 3622.787145 * Education + 4905.208294 * Female

Salary is depend on the female variable when all other conditions held constant

Salary difference between Male and Female = 4905.208294 * Female - 4905.208294 * Male

or

Salary difference = 4905.208294*(Female - Male)

But given If female =1,If male = 0

So, Salary difference = 4905.208294 * 1 - 4905.208294 * 0 = 4905.208294